OpenStudy (newwar):

Select all statements that are true about the linear equation. y=4x−3y=4x−3 4x−y=−34x−y=−3 has the same graph. The graph of the equation is a single point, representing one solution to the equation. Since the point (0, −3)(0, −3) is a solution to the equation, it is on the graph of the equation. The point (1, 1)(1, 1) is on the graph of the equation. The graph of the equation is the set of all points that are solutions to the equation.

OpenStudy (jackthegreatest):

a is true for sure, the solution of a system of equations is always a point

OpenStudy (jackthegreatest):

for b, (0,-3)(0,-3) typo?

OpenStudy (newwar):

wait A is 4x−y=−34x−y=−3 has the same graph. @jackthegreatest or did u know that

OpenStudy (jackthegreatest):

actually nvm, i am confused by the equation now

OpenStudy (jackthegreatest):

take away my medal

OpenStudy (newwar):

it's ok

OpenStudy (jackthegreatest):

haha i just realized, those equations are not simplified

OpenStudy (jackthegreatest):

no wonder they are so confusing

OpenStudy (danjs):

What is the given equation? y = 4*x - 3 hard to tell what it is

OpenStudy (newwar):

o lol

OpenStudy (jackthegreatest):

actually i am more confused now